# Countably compact inverse semigroups and Nyikos problem

### v1

@@ Abstract.
Key words and phrases:
Nyikos problem, countably compact semigroup, compact inverse semigroup, continuity of inversion
2020 Mathematics Subject Classification:
20M18, 22A15, 22A26, 54D30
1. Introduction
All topological spaces in this paper are assumed to be Hausdorff. A regular separable first-countable countably compact space is called a Nyikos space. Recall that a space $X$ is called separable if $X$ has a countable dense subset; first-countable if each point of $X$ has a countable open neighborhood base; countably compact if each infinite subset of $X$ has an accumulation point. The following problem was posed by Nyikos in 1986 and listed among 20 central problems in Set-theoretic Topology by Hrušak and Moore [33] (for other occurrences of this problem see [40, 41, 44]).
@@ 1. Introduction
Franklin and Rajagopalan [23] constructed a normal locally compact noncompact Nyikos space under the consistent assumption $\omega_{1}=\mathfrak{t}$ (for basic information about cardinal characteristics of the continuum we refer the reader to [12]). Ostaszewski [43] constructed a perfectly normal locally compact hereditary separable noncompact Nyikos space assuming $\diamondsuit$ (for more about the diamond axiom see [35, Chapter II.7]). Later van Douwen observed that Ostaszewski’s arguments yield the existence of a locally compact noncompact Nyikos space under $\mathfrak{b}=\mathfrak{c}$. Bardyla and Zdomskyy [10] constructed a consistent example of a Nyikos space which is $\mathbb{R}$-rigid, i.e. admits only constant continuous real-valued functions. Note that $\mathbb{R}$-rigid spaces, being not Tychonoff, are not locally compact. Moreover, Bardyla, Nyikos and Zdomskyy [9] showed that assuming $\omega_{1}<\mathfrak{b}=\mathfrak{s}=\mathfrak{c}$, every regular separable first-countable non-normal space of weight $<\mathfrak{c}$ embeds into an $\mathbb{R}$-rigid Nyikos space. On the other hand, Weiss [55] proved that under Martin’s Axiom every perfectly normal Nyikos space is compact. Nyikos and Zdomskyy [42] showed that the Proper Forcing Axiom, or briefly PFA, implies that each normal Nyikos space is compact. More information about PFA can be found in [38, 53]. For other fruitful applications of PFA in Topology see [17, 18, 49].
A semigroup $S$ endowed with a topology is called a topological semigroup if the semigroup operation viewed as a map from $S{\times}S$ to $S$ is continuous. For the sake of brevity, we call a topological semigroup Nyikos if its underlying space is Nyikos. Recall that each topological space possesses a compatible semigroup operation. Thus the existence of noncompact Nyikos topological semigroups is consistent with ZFC. In this paper, we consider the following natural problem:
@@
By $\omega_{1}$ we denote the first uncountable cardinal. The proof of Theorem A is based on the following two results that are of independent interest. The first one characterizes compact Nyikos topological semilattices in terms of their chains.
Theorem B.
@@
The second result solves the problem of Banakh and Pastukhova [5, Problem 2.2] and complements [5, Proposition 2.1][27, Theorem 7] and [4, Corollary 1].
Theorem C.
@@
Recall that a space $X$ is called sequential if for every non-closed subset $A\subset X$ there exists a sequence in $A$ that converges to some point of $X\setminus A$. Since each first-countable space is sequential, Theorem C implies the following.
@@
A space $X$ is called pseudocompact if $X$ is Tychonoff and each continuous real-valued function on $X$ is bounded. A semigroup $S$ endowed with a topology is called topologically periodic if for each $x\in S$ and open neighborhood $U$ of $x$ there exists $n\geq 2$ such that $x^{n}\in U$. Theorem C is essential for the proof of the following result that complements [4, Theorem 2] and [27, Theorem 1].
Theorem D.
@@
This paper is organized as follows. Section 2 is devoted to compact-like semigroups with dense inverse subsemigroups and the automatic continuity of inversion. In particular, Theorems C and D are proven there. In Section 3 we investigate chains in Nyikos topological semilattices and prove Theorem B. In Section 4 we show that each locally compact Nyikos topological semilattice is compact. This result is a milestone in the proof of Theorem A that is given in Section 5. Also, there we show that Theorem A cannot be generalized over simple bands.
2. Compact-like semigroups with dense inverse subsemigroups and the automatic continuity of inversion
@@
$s(A)=\{y\in X\colon\hbox{ there exists a sequence }\{x_{n}:n\in\omega\}\subseteq A\hbox{ such that }\lim_{n\in\omega}x_{n}=y\}.$
@@
Note that $\operatorname{cls}^{0}(A)=A$ is a subsemigroup of $S$. Assume that for each $\xi<\delta\leq\omega_{1}$, $\operatorname{cls}^{\xi}(A)$ is a semigroup. If $\delta$ is limit, then $\operatorname{cls}^{\delta}(A)=\bigcup_{\xi<\delta}\operatorname{cls}^{\xi}(A)$ is a subsemigroup of $S$, being an ascending union of semigroups. Suppose that $\delta=\gamma+1$ for some $\gamma<\omega_{1}$. Then $\operatorname{cls}^{\delta}(A)=s(\operatorname{cls}^{\gamma}(A))$. Pick any points $x,y\in\operatorname{cls}^{\delta}(A)$ and sequences $\{x_{n}:n\in\omega\}\subset\operatorname{cls}^{\gamma}(A)$ and $\{y_{n}:n\in\omega\}\subset\operatorname{cls}^{\gamma}(A)$ such that $x=\lim_{n\in\omega}x_{n}$ and $y=\lim_{n\in\omega}y_{n}$. Since $\operatorname{cls}^{\gamma}(A)$ is a semigroup, we get that the sequence $\{x_{n}y_{n}:n\in\omega\}$ is contained in $\operatorname{cls}^{\gamma}(A)$. Since $S$ is a topological semigroup, the sequence $\{x_{n}y_{n}:n\in\omega\}$ converges to $xy$. Thus $xy\in\operatorname{cls}^{\delta}(A)$, witnessing that $\operatorname{cls}^{\delta}(A)$ is a subsemigroup of $S$. Hence $\operatorname{cls}^{\alpha}(A)$ is a subsemigroup of $S$ for each $\alpha\leq\omega_{1}$, as required. ∎
The following problem appears in [5, Problem 2.2].
@@
For a semigroup $S$ by $E(S)$ we denote the set of all idempotents of $S$. If $S$ is an inverse semigroup, then $E(S)$ is a semilattice. A semigroup $S$ is called regular if for each $x\in S$ there exists $y\in S$ such that $xyx=x$. Note that every inverse semigroup is regular. A regular semigroup $S$ is inverse if an only if $E(S)$ is a semilattice (see [31, Theorem 5.1.1]). A filter $\mathcal{F}$ on a space $X$ converges to a point $x\in X$ if $\mathcal{F}$ contains all open neighborhoods of $x$. The following proposition is useful for detecting inverse semigroups among topological semigroups with a dense inverse subsemigroup.
@@
First let us show that the semigroup $S$ is regular. Fix any $y\in S\setminus X$. By the assumption, there exists an ultrafilter $\mathcal{F}$ on $S$ convergent to $y$ such that $X\in\mathcal{F}$ and the ultrafilter $\mathcal{F}^{-1}$ converges to some element $z\in S$. Consider an open neighborhood $U$ of the point $w=yzy$. By the continuity of the semigroup operation in $S$, there exist open neighborhoods $V(y)$ and $V(z)$ of $y$ and $z$, respectively, such that $V(y)V(z)V(y)\subseteq U$. Then there exist $X\supseteq F_{1}\in\mathcal{F}$ and $X\supseteq F_{2}\in\mathcal{F}$ such that $F_{1}\subset V(y)$ and $F_{2}^{-1}\subseteq V(z)$. Then for $F=F_{1}\cap F_{2}$ we get $F\subseteq FF^{-1}F\subseteq U$. Since the open neighborhood $U$ of $w$ is arbitrarily chosen, the filter $\mathcal{F}$ converges to $w$. Since the space $S$ is Hausdorff, any filter on $S$ converges to at most one point, implying that $y=w=yzy$. Hence $S$ is a regular semigroup.
@@
To prove that $S$ is an inverse semigroup it suffices to show that idempotents of $S$ commute. For this it is enough to show that $E(S)=\overline{E(X)}$, as $E(X)$ is a commutative semigroup. Fix an idempotent $e\in S\setminus X$. By the assumption, there exists an ultrafilter $\mathcal{F}$ convergent to $e$ such that $X\in\mathcal{F}$ and the ultrafilter $\mathcal{F}^{-1}$ converges to some point $f\in S$. Similarly as above one can check that $fef=f$ and $e=efe$. Since $ee=e$, the continuity of the semigroup operation implies that for every open neighborhood $U$ of $e$ there exists an open neighborhood $V$ of $e$ such that $VV\subset U$. Pick any $F\in\mathcal{F}$ such that $F\subset V$ and observe that $FF\subset U$. Consequently, the filter $\mathcal{F}^{2}$ generated by the family $\{FF:F\in\mathcal{F}\}$ converges to $e$. Fix an open neighborhood $U$ of $f$. Since $f=fef$ and $S$ is a topological semigroup, there exist open neighborhoods $V(f)$ and $V(e)$ of $f$ and $e$, respectively, such that $V(f)V(e)V(f)\subset U$. Since the filter $\mathcal{F}^{-1}$ converges to $f$ and the filter $\mathcal{F}^{2}$ converges to $e$, there exists $F\in\mathcal{F}$ such that $F\subseteq X$, $F^{-1}\subset V(f)$ and $FF\subset V(e)$. Fix any $x\in F$. Observe that
$x^{-1}xxx^{-1}\in F^{-1}FFF^{-1}\subset V(f)V(e)V(f)\subseteq U.$
@@
Since $X$ is an inverse semigroup, the element $x^{-1}xxx^{-1}$ is an idempotent. As the open neighborhood $U$ of $f$ is arbitrarily chosen, we get $f\in\overline{E(X)}$. Since $S$ is a Hausdorff topological semigroup, $f$ is an idempotent. Let $\mathcal{F}^{-2}$ be the filter generated by the family $\{FF:F\in\mathcal{F}^{-1}\}$. Similarly as above one can check that $\mathcal{F}^{-2}$ converges to $f$. Since $S$ is a topological semigroup and $e=efe$, for any open neighborhood $W$ of $e$ there exist open neighborhoods $V(f)$ and $V(e)$ of $f$ and $e$, respectively, such that $V(e)V(f)V(e)\subseteq W$. As the filter $\mathcal{F}^{-2}$ converges to $f$ and the filter $\mathcal{F}$ converges to $e$, there exists $F\in\mathcal{F}$ such that $F\subseteq X\cap V(e)$ and $(FF)^{-1}=F^{-1}F^{-1}\subset V(f)$. Then for any $x\in F$ we have
$E(X)\ni xx^{-1}x^{-1}x\in FF^{-1}F^{-1}F\subset V(e)V(f)V(e)\subseteq W.$
@@
A space $X$ is called sequentially compact if each sequence in $X$ contains a convergent subsequence. We are in a position to prove Theorem C and, in particular, to solve Problem 2.2. We need to show that a topological semigroup $S$ containing a dense inverse subsemigroup is a topological inverse semigroup provided (i) $S$ is compact, or (ii) $S$ is countably compact and sequential.
Proof of Theorem C.
(i) Let $S$ be a compact topological semigroup containing a dense inverse subsemigroup $X$. Fix any $y\in S\setminus X$ and consider an ultrafilter $\mathcal{F}$ on $S$ that contains the family $\{U\cap X:U$ is an open neighborhood of $y\}$. It is clear that $\mathcal{F}$ converges to $y$. By the compactness of $S$, the ultrafilter $\mathcal{F}^{-1}$ generated by the family $\{F^{-1}:X\supseteq F\in\mathcal{F}\}$ converges to some point of $S$. Proposition 2.3 implies that $S$ is an inverse topological semigroup. By [28, Proposition 1.6.7], inversion is automatically continuous in compact inverse topological semigroups. Hence $S$ is a topological inverse semigroup.
@@
First let us show that $S$ is an inverse semigroup. Note that $\operatorname{cls}^{0}(X)=X$ is an inverse subsemigroup of $S$. Assume that $\operatorname{cls}^{\xi}(X)$ is an inverse subsemigroup of $S$ for every $\xi<\alpha\leq\omega_{1}$. If the ordinal $\alpha$ is limit, then $\operatorname{cls}^{\alpha}(X)=\bigcup_{\xi<\alpha}\operatorname{cls}^{\xi}(X)$ is an inverse semigroup, being an ascending union of inverse semigroups. Suppose that $\alpha=\gamma+1$ for some $\gamma<\omega_{1}$. Lemma 2.1 implies that $\operatorname{cls}^{\alpha}(X)=s(\operatorname{cls}^{\gamma}(X))$ is a subsemigroup of $S$ that contains densely the inverse semigroup $\operatorname{cls}^{\gamma}(X)$. Fix any $x\in\operatorname{cls}^{\alpha}(X)\setminus\operatorname{cls}^{\gamma}(X)$ and a sequence $\{x_{n}:n\in\omega\}\subset\operatorname{cls}^{\gamma}(X)$ convergent to $x$. By [22, Theorem 3.10.31], each Hausdorff countably compact sequential space is sequentially compact. Thus the sequence $\{x_{n}^{-1}:n\in\omega\}\subseteq\operatorname{cls}^{\gamma}(X)$ contains a convergent in $S$ subsequence $\{x^{-1}_{n_{k}}:k\in\omega\}$. Set $y=\lim_{k\in\omega}x^{-1}_{n_{k}}$ and note that $y\in s(\operatorname{cls}^{\gamma}(X))=\operatorname{cls}^{\alpha}(X)$. Fix any free ultrafilter $\mathcal{F}$ on $\operatorname{cls}^{\alpha}(X)$ which contains the set $\{x_{n_{k}}:k\in\omega\}$. It is easy to see that $\mathcal{F}$ converges to $x$, and the ultrafilter $\mathcal{F}^{-1}$ generated by the family $\{F^{-1}\colon\operatorname{cls}^{\gamma}(X)\supseteq F\in\mathcal{F}\}$ converges to $y$. Proposition 2.3 implies that $\operatorname{cls}^{\alpha}(X)$ is an inverse semigroup. Hence $S=\operatorname{cls}^{\omega_{1}}(X)$ is an inverse semigroup. By [4, Corollary 1], inversion is automatically continuous on every countably compact sequential inverse topological semigroup. Hence, $S$ is a topological inverse semigroup. ∎
The bicyclic monoid is generated by two elements $p,q$ subject to the one condition $qp=1$. It is well known that bicyclic monoid is an inverse semigroup (see [36, Chapter 3.4]). The following two results that appear in [3, Theorem 6.1 and Theorem 6.6] establish some sharpness of Theorem C.
@@
Recall that an inverse semigroup $S$ is called Clifford if $xx^{-1}=x^{-1}x$ for all $x\in S$. The automatic continuity of inversion in groups, Clifford semigroups, and inverse semigroups is widely studied (see [4, 6, 14, 21, 27, 46, 48, 50]). In most cases the continuity of inversion follows from the continuity of multiplication and some compact-like property. For instance, the following two results appear in [27, Corollary 2] and [4, Theorem 2], respectively.
@@
The rest of this section is devoted to the proof of Theorem D, which in fact generalizes Theorems 2.6 and 2.7. A semigroup $S$ endowed with a topology is called semitopological if for every $s\in S$ the shifts $l_{s}:x\mapsto sx$ and $r_{s}:x\mapsto xs$ are continuous in $S$. It is clear that every topological semigroup is semitopological.
@@
Fix an element $x\in S$ and an open neighborhood $U$ of $x$. Since $X$ is topologically periodic, there exists a positive integer $n\geq 2$ such that $x^{n}\in U$. Then
@@
It follows that $V\cap(x^{2}x^{-2})V\neq\varnothing$ and $V(x^{-2}x^{2})\cap V\neq\varnothing$ for each open neighborhood $V$ of $x$. Since $X$ is a Hausdorff semitopological semigroup, we obtain $x^{2}x^{-2}x=x=xx^{-2}x^{2}$. It follows that $x^{-1}x^{2}x^{-2}=x^{-1}=x^{-2}x^{2}x^{-1}$. Then
$x^{-1}x=x^{-1}x^{2}x^{-2}x^{2}x^{-2}x=x^{-1}x^{2}x^{-2}x=(x^{-1}x)(xx^{-1})(x^{-1}x)=(xx^{-1})(x^{-1}x).$

### v2

# Countably compact inverse semigroups and Nyikos’ problem

@@ Abstract
2020 Mathematics Subject Classification: Primary 20M18, 22A15, 22A26, 54D30.Key words and phrases: Nyikos’ problem, countably compact semigroup, compact inverse semigroup, continuity of inversion.
1 Introduction
All topological spaces in this paper are assumed to be Hausdorff. A regular separable first-countable countably compact space is called a Nyikos space. Recall that a space $X$ is called separable if $X$ has a countable dense subset; first-countable if each point of $X$ has a countable neighborhood base; countably compact if each infinite subset of $X$ has an accumulation point. The following problem posed by Nyikos in 1986 is listed among 20 central problems in Set-theoretic Topology by Hrušak and Moore [34] (for other occurrences of this problem see [41, 42, 45]).
@@ 1 Introduction
Franklin and Rajagopalan [24] constructed a normal locally compact noncompact Nyikos space under the consistent assumption $\omega_{1}=\mathfrak{t}$ (for basic information about cardinal characteristics of the continuum we refer the reader to [13]). Ostaszewski [44] constructed a perfectly normal locally compact hereditary separable noncompact Nyikos space assuming $\diamondsuit$ (for more about the diamond axiom see [36, Chapter II.7]). Later van Douwen observed that Ostaszewski’s arguments yield the existence of a locally compact noncompact Nyikos space under $\mathfrak{b}=\mathfrak{c}$. Bardyla and Zdomskyy [11] constructed a consistent example of a Nyikos space which is $\mathbb{R}$-rigid, i.e. admits only constant continuous real-valued functions. Note that $\mathbb{R}$-rigid spaces, being not Tychonoff, are not locally compact. Moreover, Bardyla, Nyikos and Zdomskyy [10] showed that assuming $\omega_{1}<\mathfrak{b}=\mathfrak{s}=\mathfrak{c}$, every regular separable first-countable non-normal space of weight $\mathfrak{c}$ embeds into an $\mathbb{R}$-rigid Nyikos space. On the other hand, Weiss [57] proved that under Martin’s Axiom every perfectly normal Nyikos space is compact. Nyikos and Zdomskyy [43] showed that the Proper Forcing Axiom, or briefly PFA, implies that each normal Nyikos space is compact. More information about PFA can be found in [39, 55]. For other fruitful applications of PFA in Topology see [18, 19, 51].
A semigroup $S$ endowed with a topology is called a topological semigroup if the semigroup operation viewed as a map from $S{\times}S$ to $S$ is continuous. For the sake of brevity, we call a topological semigroup Nyikos if its underlying space is Nyikos. Each topological space $X$ possesses a continuous semigroup operation. For example, for a fixed $a\in X$ one can put $xy=a$ for all $x,y\in X$. Thus the existence of noncompact Nyikos topological semigroups is consistent with ZFC. In this paper, we consider the following natural problem:
@@
By $(\omega_{1},\min)$ and $(\omega_{1},\max)$ we denote the first uncountable cardinal $\omega_{1}$ endowed with the semilattice operations $\min$ and $\max$, respectively. The proof of Main Theorem 1.4 is based on the following two results that are of independent interest. The first one characterizes compact Nyikos topological semilattices in terms of their linear subsets (or briefly chains).
Theorem 1.5.
@@
The second result solves the problem of Banakh and Pastukhova [6, Problem 2.2] and complements [6, Proposition 2.1][28, Theorem 7] and [5, Corollary 1].
Theorem 1.6.
@@
Recall that a space $X$ is called sequential if for every non-closed subset $A\subseteq X$ there exists a sequence in $A$ that converges to some point of $X\setminus A$. Since each first-countable space is sequential, Theorem 1.6 implies the following.
@@
A space $X$ is called pseudocompact if $X$ is Tychonoff and each continuous real-valued function on $X$ is bounded. A semigroup $S$ endowed with a topology is called topologically periodic if for each $x\in S$ and neighborhood $U$ of $x$ there exists $n\geq 2$ such that $x^{n}\in U$. Theorem 1.6 is essential for the proof of the following result that complements [5, Theorem 2] and [28, Theorem 1].
Theorem 1.8.
@@
This paper is organized as follows. Section 2 is devoted to compact-like semigroups with dense inverse subsemigroups and the automatic continuity of inversion. In particular, Theorems 1.6 and 1.8 are proven there. In Section 3 we investigate chains in Nyikos topological semilattices and prove Theorem 1.5. In Section 4 we show that each locally compact Nyikos topological semilattice is compact. This result is a milestone in the proof of Main Theorem 1.4 that is given in Section 5. Also, there we show that Main Theorem 1.4 cannot be generalized over simple bands.
2 Compact-like semigroups with dense inverse subsemigroups and the automatic continuity of inversion
@@
$s(A)=\{y\in X\colon\hbox{ }\exists\{x_{n}:n\in\omega\}\subseteq A\hbox{ such that }\lim_{n\in\omega}x_{n}=y\}.$
@@
Note that $\operatorname{cls}^{0}(A)=A$ is a subsemigroup of $S$. Assume that for each $\xi<\delta\leq\omega_{1}$, $\operatorname{cls}^{\xi}(A)$ is a semigroup. If $\delta$ is limit, then $\operatorname{cls}^{\delta}(A)=\bigcup_{\xi<\delta}\operatorname{cls}^{\xi}(A)$ is a subsemigroup of $S$, being an ascending union of semigroups. Suppose that $\delta=\gamma+1$ for some $\gamma<\omega_{1}$. Then $\operatorname{cls}^{\delta}(A)=s(\operatorname{cls}^{\gamma}(A))$. Pick any points $x,y\in\operatorname{cls}^{\delta}(A)$ and sequences $\{x_{n}:n\in\omega\}\subseteq\operatorname{cls}^{\gamma}(A)$ and $\{y_{n}:n\in\omega\}\subseteq\operatorname{cls}^{\gamma}(A)$ such that $x=\lim_{n\in\omega}x_{n}$ and $y=\lim_{n\in\omega}y_{n}$. Since $\operatorname{cls}^{\gamma}(A)$ is a semigroup, we get that the sequence $\{x_{n}y_{n}:n\in\omega\}$ is contained in $\operatorname{cls}^{\gamma}(A)$. Since $S$ is a topological semigroup, the sequence $\{x_{n}y_{n}:n\in\omega\}$ converges to $xy$. Thus $xy\in\operatorname{cls}^{\delta}(A)$, witnessing that $\operatorname{cls}^{\delta}(A)$ is a subsemigroup of $S$. Hence $\operatorname{cls}^{\alpha}(A)$ is a subsemigroup of $S$ for each $\alpha\leq\omega_{1}$, as required. ∎
The following problem appears in [6, Problem 2.2].
@@
For a semigroup $S$ by $E(S)$ we denote the set of all idempotents of $S$. A semigroup $S$ is called regular if for each $x\in S$ there exists $y\in S$ such that $xyx=x$. Note that every inverse semigroup is regular. A regular semigroup $S$ is inverse if an only if $E(S)$ is a semilattice (see [32, Theorem 5.1.1]). A filter $\mathcal{F}$ on a space $X$ converges to a point $x\in X$ if $\mathcal{F}$ contains all neighborhoods of $x$. For a subset $F$ of an inverse semigroup $S$ let $F^{-1}=\{x^{-1}:x\in F\}$. The following proposition is useful for detecting inverse semigroups among topological semigroups with a dense inverse subsemigroup.
@@
First let us show that the semigroup $S$ is regular. Fix any $y\in S\setminus X$. By assumption, there exists an ultrafilter $\mathcal{F}$ on $S$ convergent to $y$ such that $X\in\mathcal{F}$ and the ultrafilter $\mathcal{F}^{-1}$ converges to some element $z\in S$. Consider a neighborhood $U$ of the point $w=yzy$. By the continuity of the semigroup operation in $S$, there exist neighborhoods $V(y)$ and $V(z)$ of $y$ and $z$, respectively, such that $V(y)V(z)V(y)\subseteq U$. Then there exist $F_{1},F_{2}\in\mathcal{F}$ such that $F_{1}\subseteq V(y)\cap X$ and $F_{2}^{-1}\subseteq V(z)\cap X$. Then for $F=F_{1}\cap F_{2}$ we get $F\subseteq FF^{-1}F\subseteq U$. Since the neighborhood $U$ of $w$ is arbitrarily chosen, the filter $\mathcal{F}$ converges to $w$. Since the space $S$ is Hausdorff, any filter on $S$ converges to at most one point, implying that $y=w=yzy$. Hence $S$ is a regular semigroup.
@@
To prove that $S$ is an inverse semigroup it suffices to show that idempotents of $S$ commute. For this it is enough to show that $E(S)=\overline{E(X)}$, as $E(X)$ is a commutative semigroup. Fix an idempotent $e\in S\setminus X$. By assumption, there exists an ultrafilter $\mathcal{F}$ convergent to $e$ such that $X\in\mathcal{F}$ and the ultrafilter $\mathcal{F}^{-1}$ converges to some point $f\in S$. Similarly as above one can check that $fef=f$ and $e=efe$. Since $ee=e$, the continuity of the semigroup operation implies that for every neighborhood $U$ of $e$ there exists a neighborhood $V$ of $e$ such that $VV\subseteq U$. Pick any $F\in\mathcal{F}$ such that $F\subseteq V$ and observe that $FF\subseteq U$. Consequently, the filter $\mathcal{F}^{2}$ generated by the family $\{FF:F\in\mathcal{F}\}$ converges to $e$. Fix a neighborhood $U$ of $f$. Since $f=fef$ and $S$ is a topological semigroup, there exist neighborhoods $V(f)$ and $V(e)$ of $f$ and $e$, respectively, such that $V(f)V(e)V(f)\subseteq U$. Since the filter $\mathcal{F}^{-1}$ converges to $f$ and the filter $\mathcal{F}^{2}$ converges to $e$, there exists $F\in\mathcal{F}$ such that $F\subseteq X$, $F^{-1}\subseteq V(f)$ and $FF\subseteq V(e)$. Fix any $x\in F$. Observe that
$x^{-1}xxx^{-1}\in F^{-1}FFF^{-1}\subseteq V(f)V(e)V(f)\subseteq U.$
@@
Since $X$ is an inverse semigroup, the elements $x^{-1}x$ and $xx^{-1}$ are idempotents. Since $E(X)$ is a subsemigroup of $X$, the element $x^{-1}xxx^{-1}$ is also an idempotent. As the neighborhood $U$ of $f$ is arbitrarily chosen, we get $f\in\overline{E(X)}$. Since $S$ is a Hausdorff topological semigroup, $f$ is an idempotent. Let $\mathcal{F}^{-2}$ be the filter generated by the family $\{FF:F\in\mathcal{F}^{-1}\}$. Similarly as above one can check that $\mathcal{F}^{-2}$ converges to $f$. Since $S$ is a topological semigroup and $e=efe$, for any neighborhood $W$ of $e$ there exist neighborhoods $V(f)$ and $V(e)$ of $f$ and $e$, respectively, such that $V(e)V(f)V(e)\subseteq W$. As the filter $\mathcal{F}^{-2}$ converges to $f$ and the filter $\mathcal{F}$ converges to $e$, there exists $F\in\mathcal{F}$ such that $F\subseteq X\cap V(e)$ and $(FF)^{-1}=F^{-1}F^{-1}\subseteq V(f)$. Then for any $x\in F$ we have
$E(X)\ni xx^{-1}x^{-1}x\in FF^{-1}F^{-1}F\subseteq V(e)V(f)V(e)\subseteq W.$
@@
A space $X$ is called sequentially compact if each sequence in $X$ contains a convergent subsequence. We are in a position to prove Theorem 1.6 and, in particular, to solve Problem 2.2. We need to show that a topological semigroup $S$ containing a dense inverse subsemigroup is a topological inverse semigroup provided (i) $S$ is compact, or (ii) $S$ is countably compact and sequential.
Proof of Theorem 1.6.
(i) Let $S$ be a compact topological semigroup containing a dense inverse subsemigroup $X$. Fix any $y\in S\setminus X$ and consider an ultrafilter $\mathcal{F}$ on $S$ that contains the family $\{U\cap X:U$ is a neighborhood of y}y\}. It is clear that $\mathcal{F}$ converges to $y$. By the compactness of $S$, the ultrafilter $\mathcal{F}^{-1}$ generated by the family $\{F^{-1}:X\supseteq F\in\mathcal{F}\}$ converges to some point of $S$. Proposition 2.3 implies that $S$ is an inverse topological semigroup. By [29, Proposition 1.6.7], inversion is automatically continuous in compact inverse topological semigroups. Hence $S$ is a topological inverse semigroup.
@@
First let us show that $S$ is an inverse semigroup. Note that $\operatorname{cls}^{0}(X)=X$ is an inverse subsemigroup of $S$. Assume that $\operatorname{cls}^{\xi}(X)$ is an inverse subsemigroup of $S$ for every $\xi<\alpha\leq\omega_{1}$. If the ordinal $\alpha$ is limit, then $\operatorname{cls}^{\alpha}(X)=\bigcup_{\xi<\alpha}\operatorname{cls}^{\xi}(X)$ is an inverse semigroup, being an ascending union of inverse semigroups. Suppose that $\alpha=\gamma+1$ for some $\gamma<\omega_{1}$. Lemma 2.1 implies that $\operatorname{cls}^{\alpha}(X)=s(\operatorname{cls}^{\gamma}(X))$ is a subsemigroup of $S$ that contains densely the inverse semigroup $\operatorname{cls}^{\gamma}(X)$. Fix any $x\in\operatorname{cls}^{\alpha}(X)\setminus\operatorname{cls}^{\gamma}(X)$ and a sequence $\{x_{n}:n\in\omega\}\subseteq\operatorname{cls}^{\gamma}(X)$ convergent to $x$. By [23, Theorem 3.10.31], each Hausdorff countably compact sequential space is sequentially compact. Thus the sequence $\{x_{n}^{-1}:n\in\omega\}\subseteq\operatorname{cls}^{\gamma}(X)$ contains a convergent in $S$ subsequence $\{x^{-1}_{n_{k}}:k\in\omega\}$. Set $y=\lim_{k\in\omega}x^{-1}_{n_{k}}$ and note that $y\in s(\operatorname{cls}^{\gamma}(X))=\operatorname{cls}^{\alpha}(X)$. Fix any free ultrafilter $\mathcal{F}$ on $\operatorname{cls}^{\alpha}(X)$ which contains the set $\{x_{n_{k}}:k\in\omega\}$. It is easy to see that $\mathcal{F}$ converges to $x$, and the ultrafilter $\mathcal{F}^{-1}$ generated by the family $\{F^{-1}\colon\operatorname{cls}^{\gamma}(X)\supseteq F\in\mathcal{F}\}$ converges to $y$. Proposition 2.3 implies that $\operatorname{cls}^{\alpha}(X)$ is an inverse semigroup. Hence $S=\operatorname{cls}^{\omega_{1}}(X)$ is an inverse semigroup. By [5, Corollary 1], inversion is automatically continuous on every countably compact sequential inverse topological semigroup. Hence, $S$ is a topological inverse semigroup. ∎
The bicyclic monoid is generated by two elements $p,q$ subject to one condition $qp=1$. It is well known that the bicyclic monoid is an inverse semigroup (see [37, Chapter 3.4]). The following two results that appear in [4, Theorem 6.1 and Theorem 6.6] establish some sharpness of Theorem 1.6.
@@
Recall that an inverse semigroup $S$ is called Clifford if $xx^{-1}=x^{-1}x$ for all $x\in S$. The automatic continuity of inversion in groups, Clifford semigroups, and inverse semigroups is widely studied (see [5, 7, 15, 22, 28, 47, 48, 50, 52]). In most cases the continuity of inversion follows from the continuity of multiplication and some compact-like property. For instance, the following two results appear in [28, Corollary 2] and [5, Theorem 2], respectively.
@@
The rest of this section is devoted to the proof of Theorem 1.8, which in fact generalizes Theorems 2.6 and 2.7. A semigroup $S$ endowed with a topology is called semitopological if for every $s\in S$ the shifts $l_{s}:x\mapsto sx$ and $r_{s}:x\mapsto xs$ are continuous in $S$. It is clear that every topological semigroup is semitopological.
@@
Fix an element $x\in S$ and a neighborhood $U$ of $x$. Since $X$ is topologically periodic, there exists a positive integer $n\geq 2$ such that $x^{n}\in U$. Then
@@
It follows that $V\cap(x^{2}x^{-2})V\neq\varnothing$ and $V(x^{-2}x^{2})\cap V\neq\varnothing$ for each neighborhood $V$ of $x$. Since $X$ is a Hausdorff semitopological semigroup, we obtain $x^{2}x^{-2}x=x=xx^{-2}x^{2}$. It follows that $x^{-1}x^{2}x^{-2}=x^{-1}=x^{-2}x^{2}x^{-1}$. Then
$x^{-1}x=x^{-1}x^{2}x^{-2}x^{2}x^{-2}x$ $=x^{-1}x^{2}x^{-2}x=$
$=(x^{-1}x)(xx^{-1})(x^{-1}x)=(xx^{-1})(x^{-1}x).$